# Mixture of Parrots [Uncaptioned image][Uncaptioned image][Uncaptioned image]: Experts improve memorization more than reasoning

### v1

@@ C.2 Memorization Proofs
In our case, let $X_{i}=\bm{1}\{\text{sample }i\text{ is routed to expert }j\}$, for some fixed $j\in[K]$. Note that $\operatorname*{\mathbb{P}}[X_{i}=1]=1/K$ due to Lemma C.6 and as such $E[S_{N}]=N/K$. Further, $X_{i}\in[0,1]$. Hence we obtain:
@@
Lemma C.8.
Fix $\delta\in(0,1)$, $K\in\mathbb{N}$ and the routing function as described in Lemma C.6. Given $n\geq\frac{K^{2}\ln\left(1/\delta\right)}{2}$ samples with embedding dimension $m$. For every expert $i$, with probability at least $1-\delta$, there exists an MLP of width $\tilde{O}(n/mK)$ that correctly classifies the examples routed to this expert.
@@
Next, we fix the number of experts to be $K$. We fix a router function $r:\mathbb{R}^{\log K}\rightarrow[K]$ such that, on average, each expert must memorize $n/K$ sequences, and with high probability, each expert must memorize at most $2n/K$ sequences. We use the construction from Lemma C.6 where the router only acts on the first $\log K$ entries of each vector. Hence, we have $K$ MLPs, each tasked with memorizing at most $\frac{2n}{K}$ input-output values. We use the network construction from Daniely (2020) to only the last $m-\log K\geq m/2$ coordinates, which are not used by the router, so that their distribution is independent and Gaussian.
@@
Lemma C.10 (Bound on $\ell_{2}$-norm of vector from $\mathcal{N}(0,cI_{m})$ ).
@@
Lemma C.11 (Bounded bit complexity required).
@@
We show that the learning result from Daniely (2020) gives a network with bounded bit complexity. Hence, it suffices to bound the bit complexity of the initial weights and the bit complexity of the gradient update. First, we begin with $h(x)=\sqrt{\frac{1}{q}}\sum_{i=1}^{q}a_{i}\sigma(\langle w_{i},x\rangle)$. The objective is to show there exists $\tilde{h}(x)$ with bounded bit complexity which satisfies $y_{i}\tilde{h}(x_{i})>0\ \forall i$. Suppose we begin by creating bins for the weights, and replace each $w_{i}$ with $\tilde{w}_{i}$ where $\|w_{i}-\tilde{w}_{i}\|_{\infty}\leq\varepsilon$. Further, assume that $\|w_{i}\|_{\infty}\leq M$ for some $M>0$. We then obtain that $\|w_{i}-\tilde{w}_{i}\|_{2}\leq\sqrt{m}\varepsilon$. In addition, we state that with probability at least $1-\delta$, $\|x\|_{2}\leq c_{\delta,m}$ due to Lemma C.10. Using this, we have that:
@@
Consider $w_{i}^{(0)}$ to be the initialization of $w_{i}$ prior to the gradient step. Then, because, $M=\|w_{i}\|_{\infty}$, we have that $M=\max_{i}\left\{{w_{i}^{(0)}}-\eta\frac{\partial h(x)}{\partial w_{i}}\right\}=\max_{i}\left\{{w_{i}^{(0)}}-\frac{n\log m}{m}\frac{\partial h(x)}{\partial w_{i}}\right\}$. Assuming that $w_{i}^{(0)}$ is initialized with bounded $\ell_{\infty}$ norm of $1$, then we obtain that $M\leq 1+\frac{n\log m}{m\sqrt{q}}c_{\delta,m}$. This is due to Lemma C.10 and that
@@
Theorem C.12 (Lower bound for dense model).
@@
Lemma C.13 (Active parameter comparison between dense and sparse transformers).

### v2

@@ D.2 Memorization Proofs
In our case, let $X_{i}=\bm{1}\{\text{sample }i\text{ is routed to expert }j\}$, for some fixed $j\in[K]$. Note that $\operatorname*{\mathbb{P}}[X_{i}=1]=1/K$ due to Lemma D.6 and as such $E[S_{N}]=N/K$. Further, $X_{i}\in[0,1]$. Hence we obtain:
@@
Lemma D.8.
Fix $\delta\in(0,1)$, $K\in\mathbb{N}$ and the routing function as described in Lemma D.6. Given $n\geq\frac{K^{2}\ln\left(1/\delta\right)}{2}$ samples with embedding dimension $m$. For every expert $i$, with probability at least $1-\delta$, there exists an MLP of width $\tilde{O}(n/mK)$ that correctly classifies the examples routed to this expert.
@@
Next, we fix the number of experts to be $K$. We fix a router function $r:\mathbb{R}^{\log K}\rightarrow[K]$ such that, on average, each expert must memorize $n/K$ sequences, and with high probability, each expert must memorize at most $2n/K$ sequences. We use the construction from Lemma D.6 where the router only acts on the first $\log K$ entries of each vector. Hence, we have $K$ MLPs, each tasked with memorizing at most $\frac{2n}{K}$ input-output values. We use the network construction from Daniely (2020) to only the last $m-\log K\geq m/2$ coordinates, which are not used by the router, so that their distribution is independent and Gaussian.
@@
Lemma D.10 (Bound on $\ell_{2}$-norm of vector from $\mathcal{N}(0,cI_{m})$ ).
@@
Lemma D.11 (Bounded bit complexity required).
@@
We show that the learning result from Daniely (2020) gives a network with bounded bit complexity. Hence, it suffices to bound the bit complexity of the initial weights and the bit complexity of the gradient update. First, we begin with $h(x)=\sqrt{\frac{1}{q}}\sum_{i=1}^{q}a_{i}\sigma(\langle w_{i},x\rangle)$. The objective is to show there exists $\tilde{h}(x)$ with bounded bit complexity which satisfies $y_{i}\tilde{h}(x_{i})>0\ \forall i$. Suppose we begin by creating bins for the weights, and replace each $w_{i}$ with $\tilde{w}_{i}$ where $\|w_{i}-\tilde{w}_{i}\|_{\infty}\leq\varepsilon$. Further, assume that $\|w_{i}\|_{\infty}\leq M$ for some $M>0$. We then obtain that $\|w_{i}-\tilde{w}_{i}\|_{2}\leq\sqrt{m}\varepsilon$. In addition, we state that with probability at least $1-\delta$, $\|x\|_{2}\leq c_{\delta,m}$ due to Lemma D.10. Using this, we have that:
@@
Consider $w_{i}^{(0)}$ to be the initialization of $w_{i}$ prior to the gradient step. Then, because, $M=\|w_{i}\|_{\infty}$, we have that $M=\max_{i}\left\{{w_{i}^{(0)}}-\eta\frac{\partial h(x)}{\partial w_{i}}\right\}=\max_{i}\left\{{w_{i}^{(0)}}-\frac{n\log m}{m}\frac{\partial h(x)}{\partial w_{i}}\right\}$. Assuming that $w_{i}^{(0)}$ is initialized with bounded $\ell_{\infty}$ norm of $1$, then we obtain that $M\leq 1+\frac{n\log m}{m\sqrt{q}}c_{\delta,m}$. This is due to Lemma D.10 and that
@@
Theorem D.12 (Lower bound for dense model).
@@
Lemma D.13 (Active parameter comparison between dense and sparse transformers).